# Conformal geodesics are not variational in higher dimensions

### v1

@@ 4. Reduction and Proof in the general case
Note however that nondegeneracy condition (14) for $L$ fails in odd dimensions. Indeed, the determinant in (14) is the square of the Pfaffain of the skew-symmetric matrix $\bigl{[}\tfrac{\partial EL_{j}}{\partial x^{i}_{2}}\bigr{]}$ and in odd dimensions this always vanishes. Thus one can make a reduction by two dimensions, when restriction yields nonzero Pfaffian, and complete the induction step based on non-variationality of the parametrized conformal circles in dimensions $n=3,4$ (in fact, one could start at $n=2$ but this is not conformal but Möbius geometry):
@@
Since investigation of the equation for un-parametrized conformal circles in 5D was only briefly discussed there, let us present an important modification. Nondegeneracy condition (15) signifies equivalence of the Euler-Langrange equations to (2). Yet, one may consider overdetermined system $\mathop{\rm ev}\nolimits_{\mathcal{E}}EL_{j}$ split by $x^{i}_{2}$ even without this inequality. An important step in simplification was change by divergence $L\mapsto L+\tfrac{d}{dt}\psi$, but this can be done also when the order of $L$ is below 3, that is, without restriction to $\mathcal{E}$. In this case the Euler-Lagrange equation holds iff $L$ is a pure divergence, so it reduces to 0.
@@
3. General non-flat case. The above results show that conformal geodesics in general dimension are not variational. Let us first deduce the negative results for generic metrics $g$ from computations for flat metrics $g_{0}$. We omit large explicit formulae for equations, referring to the simplest representatives in the flat case, as this is not essential for our non-computational argument.
We adapt the approach from the discussion in the previous subsection for un-parametrized case $n=5$, without using open condition (15); the variationality is given by a system of closed differential-algebraic relations on the entry of the metric $g$ representing the conformal structure. Since these conditions fail for flat metric $ds^{2}_{\mathop{\rm Eucl}\nolimits}$ they also fail for metrics $g$ in an open neighborhood in some $C^{k}$ topology (actually, $k=5$ suffices, but we will not use it). Moreover, due to algebraicity of the conditions, the variationality fails for metrics $g$ in an open set in Zariski topology on $k$-jets of the metrics, whence for generic metrics.
In fact, we can obtain obstructions to variationality without any explicit integration of particular constraints, as in Section 3. The condition that $\delta_{X}L=0$ (without restriction to equation $\mathcal{E}$ of conformal geodesics) is equivalent to the condition of complete divergence $L\,dt=\hat{d}\psi$, where $\hat{d}$ is the horizontal differential (see [1, 21] for details on this via variational bicomplex/$\mathcal{C}$-spectral sequence).
Denote by $\Delta$ the equation of complete divergence, and by $\Delta^{\mathcal{E}}$ the Euler-Lagrange condition on equation that we previously denoted $\mathop{\rm ev}\nolimits_{\mathcal{E}}EL$. Both are submanifolds obtained by prolongations in the corresponding jet spaces (cf. [21] for details on jets and differential equations) and nested so:
$\Delta_{k}\subset\Delta_{k}^{\mathcal{E}}\subset J^{k}.$ (16)
Both $\Delta_{k},\Delta_{k}^{\mathcal{E}}$ are differential equations on components $f_{i}\in C^{\infty}(J^{1}(M^{n},1))$ of the Lagrangian $L=f_{0}+\sum_{j=1}^{n}f_{j}x_{2}^{j}$, depending on 1-jets of curves in $M$, hence the space in the right-hand-side of (16) is the space of $k$-jets of functions on $J^{1}(M^{n},1)$. Since PDEs in (16) are linear, the quotient of the corresponding jet-subbundles $\Delta_{k}^{\mathcal{E}}/\Delta_{k}$ is a bundle over $M$ co-filtered by $k$. While equation $\Delta_{k}^{\mathcal{E}}$ is of infinite type, the corresponding quotient equation on $[f_{i}]$ is of finite type. Indeed, this property we have observed for flat metrics and it does not change for general metrics, as it only depends on the symbol of the equation.
Actually, the equation $\Delta^{\mathcal{E}}$ has order 2 in $f_{i}$ and also order 2 in parameter metric $g$ (it enters through Christoffel symbols in $f_{i}$). Its third prolongation ($k=3+2$) allows to express all normal jets $[j^{k}f_{i}]\in\Delta_{k}^{\mathcal{E}}/\Delta_{k}$. In other words, the matrix $B_{g}$ of the linear system on these normal jets is of maximal rank for $g=g_{0}$ being the flat metric. But then it is also of maximal rank for any metric $g$ with generic $k$-jet, whence $[j^{k}f_{i}]\in\Delta_{k}$ and this means that $L$ is a complete divergence.
@@
Alternatively this argument can be presented as follows. The condition that a (nonlinear) differential operator $F$ is the Euler-Lagrange operator of some $L$ is given by the self-adjointness condition (here $\ell_{F}$ is the operator of universal linearization of $F$ and $\ell_{F}^{*}$ is its formal adjoint [20])
@@
The conformal geodesic equation equation $\mathcal{E}:F=0$ is equivalent to $\Box F$ for an invertible $C$-differential operator $\Box$ (we refer to [21] for the formalism); in our case the coefficients of $\Box$ can be assumed functions on $J^{1}(M,1)$. Solvability of the differential equation $\ell_{\Box F}^{*}=\ell_{\Box F}$ on these coefficients is a finite type system, which is not solvable for the parameter $g_{0}$ and hence not solvable for a generic metric $g$.
@@ 5. Discussion: on the selection principle
Conformal geodesics extend naturally to (and through) conformal infinity. Their behavior at the conformal infinity uncovers the structure of asymptotically simple solutions of the Einstein field equations [18, 9]. In infinity this family of curves carries only a projective family of parametrizations, which makes un-parametrized conformal geodesics at the three-dimensional spacial boundary physically meaningful.
Our main result states that the variationality of conformal geodesics is the selection principle for physical dimension three. In all other dimensions (including dimension two and parametrized version in dimension three) the equation is not variational. Here we test only the first variation of conformal geodesics. Indeed, even in the flat case, there exist infinitely many conformal geodesics through any two (sufficiently close) points and thus the second variation does not seem relevant here.
@@
$\mathcal{E}_{\vartriangle}\subset\mathcal{E}\subset\mathcal{E}^{\triangledown}.$
For instance, in the case of conformal geodesic equation $\mathcal{E}$ a nondegenerate conformal invariant Lagrangian of order 2 was proposed in [3], which up to divergence coincides with the one investigated further in [8]. It was noted in [3] that its Euler-Lagrange equation is of forth order; this equation $\mathcal{E}^{\triangledown}\supset\mathcal{E}$ containing the equation of conformal geodesics was later studied in [19, 16]. An example of smaller equation $\mathcal{E}_{\vartriangle}$ (for Kähler mertric $g$) is given by Kähler magnetic geodesics discussed in [8].
@@
Let the equation be given by =$\mathcal{E}=\{F_{i}=0\}_{i=1}^{m}$, where $F_{i}$ are (nonlinear) differential operators on $u=(u^{j})_{j=1}^{n}$ (usually, for determined systems, $m=n$). Then $L^{\triangledown}=\sum_{i=1}^{m}v^{i}F_{i}$ for additional dependent variable $v=(v^{i})_{i=1}^{m}$. Indeed, the equations $\delta_{v^{i}}L^{\triangledown}=F_{i}=0$ for $1\leq i\leq n$ imply $\mathcal{E}$. The other Euler-Lagrange equations $\delta_{u^{j}}L^{\triangledown}=\sum_{i=1}^{n}v^{i}\delta_{u^{j}}F_{i}=0$ should be considered as constraints on $v$. Taken together these two sets of equations define what is known as cotangent covering $\mathcal{T}^{*}\mathcal{E}$ over $\mathcal{E}$, cf. [21]. Its linearization is given by the operator $\ell_{\mathcal{E}}\oplus\ell_{\mathcal{E}}^{*}$ that is self-adjoint and hence (see e.g. [1, 20]) $\mathcal{T}^{*}\mathcal{E}$ is a variational equation. ∎
Similarly, any differential equation equation $\mathcal{E}$ (with more than one-dimensional solution space) contains a reduction to scalar second order ODE (path geometry) that is variational [10]. In other words, under mild conditions, every equation $\mathcal{E}$ contains a variational intermediate integral $\mathcal{E}_{\vartriangle}$.
Our final remark concerns variations with prescribed boundary conditions. With the standard approach one usually assumes fixed endpoint and derivatives of variations at it. Non-standard variational principle for conformal geodesics in [8] restricts the variations $V$ along the curve by $V^{\prime}(t_{0})=0$, $V^{\prime}(t_{1})=0$ as well as the boundary condition $B(V)|_{t_{0}}^{t_{1}}=-K(V)$, where $B$ is a certain second order differential operator. For the flat metric $K=0$. Consider, for instance, a one-parametric family of conformal circles in ${\mathbb{R}}^{n}$, with center $(0,\tfrac{1-c^{2}}{2c},0,\dots,0)$ and radius $\tfrac{c^{2}+1}{2c}$, passing through the points (±$(\pm 1,0,\dots,0)$ for $t=\pm 1$:
@@
.$\bigl{.}B(V)\bigr{|}_{t=-1}^{t=+1}=\frac{c^{2}-1}{c^{2}+1}\bigl{(}V_{1}^{\prime\prime}(1)-V_{1}^{\prime\prime}(-1)\bigr{)}-\frac{2c}{c^{2}+1}\bigl{(}V_{2}^{\prime\prime}(1)+V_{2}^{\prime\prime}(-1)\bigr{)}-\frac{2}{c^{2}+1}\bigl{(}V_{1}(1)-V_{1}(-1)\bigr{)}.$
@@
Acknowledgment. I would like to thank Vladimir Matveev, Wijnand Steneker, Eugene Ferapontov, Iosif Krasilshchik and Konstantin Druzhkov for useful discussions. The research was supported by the Tromsø Research Foundation (project “Pure Mathematics in Norway”) and the UiT Aurora project MASCOT.

### v2

@@ 4. Reduction and Proof in the general case
Note however that nondegeneracy condition (15) for $L$ fails in odd dimensions. Indeed, the determinant in (15) is the square of the Pfaffian of the skew-symmetric matrix $\bigl[\tfrac{\partial EL_{j}}{\partial x^{i}_{2}}\bigr]$ and in odd dimensions this always vanishes. Thus one can make a reduction by two dimensions, when restriction yields nonzero Pfaffian, and complete the induction step based on non-variationality of the parametrized conformal circles in dimensions $n=3,4$ (in fact, one could start at $n=2$ but this is not conformal but Möbius geometry):
@@
Since investigation of the equation for un-parametrized conformal circles in 5D was only briefly discussed there, let us present an important modification. Nondegeneracy condition (16) signifies equivalence of the Euler-Lagrange equations to (2). Yet, one may consider overdetermined system $\mathop{\rm ev}\nolimits_{\mathcal{E}}EL_{j}$ split by $x^{i}_{2}$ even without this inequality. An important step in simplification was change by divergence $L\mapsto L+\tfrac{d}{dt}\psi$, but this can be done also when the order of $L$ is below 3, that is, without restriction to $\mathcal{E}$. In this case the Euler-Lagrange equation holds iff $L$ is a pure divergence, so it reduces to 0.
@@
Let us make a remark on the above reduction procedure. We prove non-variationality via reduction of dimension by two in both cases: for parametrized conformal geodesics dimensions drop $2n\mapsto 2n-2\mapsto\dots$, for un-parametrized conformal geodesics dimensions drop $2n+1\mapsto 2n-1\mapsto\dots$ to keep nondegeneracy. One may expect a dimension reduction by one through switching between the two cases. This is indeed possible for the corresponding equations: for parametrized curves forget about parametrization and then restrict to a hyperplane; for unparametrized curves choose the projective parametrer and then restrict to a hyperplane. However in terms of Lagrangians the cases are distinguished by the property of the action $\int L\,dt$ being reparemetrization invariant and this does not change under restriction (of the Lagrangian density and variations) to the equation of hypersurface. Thus reduction by one dimension seems not possible.
3. General non-flat case. The above results imply that conformal geodesics in general dimension are not variational, as we shall now demonstrate. Let us first deduce the negative results for generic metrics $g$ from computations for flat metrics $g_{0}$. By generic we mean such $g$ that its $k$-jets is generic at a reference point for specified $k$. We omit large explicit formulae for equations, referring to the simplest representatives in the flat case, as this is not essential for our non-computational argument.
We adapt the approach from the discussion in the previous subsection for un-parametrized case, without using nondegeneracy condition (16); the variationality is given by a system of closed differential-algebraic relations on the entry of the metric $g$ representing the conformal structure. Since these conditions fail for flat metric $ds^{2}_{\mathop{\rm Eucl}\nolimits}$ they also fail for metrics $g$ in an open neighborhood in some $C^{k}$ topology (actually, $k=5$ suffices, but we will not use it). Moreover, due to algebraicity of the conditions, the variationality fails for metrics $g$ in an open set in Zariski topology on $k$-jets of the metrics, whence for generic metrics.
@@
Denote by $\Delta^{0}$ the equation of complete divergence, and by $\Delta^{\mathcal{E}}$ the Euler-Lagrange condition on equation that we previously denoted $\mathop{\rm ev}\nolimits_{\mathcal{E}}EL$. Both are submanifolds obtained by prolongations in the corresponding jet spaces (cf. [25] for details on jets and differential equations) and nested so:
$\Delta_{k}^{0}\subset\Delta_{k}^{\mathcal{E}}\subset J^{k}.$ (17)
Both $\Delta_{k}^{0},\Delta_{k}^{\mathcal{E}}$ are differential equations on components $f_{i}\in C^{\infty}(J^{1}(M^{n},1))$ of the Lagrangian $L=f_{0}+\sum_{j=1}^{n}f_{j}x_{2}^{j}$, depending on 1-jets of curves in $M$, hence the space in the right-hand-side of (17) is the space of $k$-jets of functions $f_{i}$ on $J^{1}(M^{n},1)$. Note that $\Delta^{0}$ expresses triviality of the Euler-Lagrange equation for the above $L$, while $\Delta^{\mathcal{E}}$ expresses this triviality modulo $\mathcal{E}$.
Since PDEs in (17) are linear, the quotient of the corresponding jet-subbundles $\Delta_{k}^{\mathcal{E}}/\Delta_{k}$ is a bundle over $M$ co-filtered by $k$. While equation $\Delta_{k}^{\mathcal{E}}$ is of infinite type, the corresponding quotient equation on $[f_{i}]$ is of finite type. It means that there are only finitely many linearly independent jets of $f_{i}$ in the quotient $\Delta^{\mathcal{E}}/\Delta^{0}$, and we call normal jets a basis of those. Indeed, this finite type property we have observed for flat metrics and it does not change for general metrics, as it only depends on the symbol of the equation.
Actually, the equation $\Delta^{\mathcal{E}}$ has order 2 in $f_{i}$ and also order 2 in parameter metric $g$ (it enters through Christoffel symbols, Schouten tensor, etc). Its third prolongation ($k=3+2$) allows to express all normal jets $[j^{k}f_{i}]\in\Delta_{k}^{\mathcal{E}}/\Delta_{k}^{0}$. In other words, the matrix $B_{g}$ of the linear system on these normal jets is of maximal rank for $g=g_{0}$ being the flat metric. But then it is also of maximal rank for any metric $g$ with generic $k$-jet, whence $[j^{k}f_{i}]\in\Delta_{k}^{0}$ and this means that $L$ is a complete divergence. Thus we have proved:
@@
Let us make a remark on the usage of a reference point (where the $k$-jet is based) at which we evaluate the rank of $B_{g}$ in the above proof. For (conformally) flat metric all points play the same role due to homogenuity. In general, there can be regular and singular points (with respect to stability of ranks of various tensors). The former are generic and we may always choose the reference point regular. However we may also take any singular point, because the argument is based on comparisons of $B_{g}$ with $B_{g_{0}}$.
@@
Alternatively the above argument can be presented as follows. The condition that a (nonlinear) differential operator $F$ is the Euler-Lagrange operator of some $L$ is given by the self-adjointness condition (here $\ell_{F}$ is the operator of universal linearization of $F$ and $\ell_{F}^{*}$ is its formal adjoint [24])
@@
The conformal geodesic equation $\mathcal{E}:F=0$ is equivalent to $\Box F$ for an invertible $C$-differential operator $\Box$ (we refer to [25] for the formalism); in our case the coefficients of $\Box$ can be assumed functions on $J^{1}(M,1)$. Solvability of the differential equation $\ell_{\Box F}^{*}=\ell_{\Box F}$ on these coefficients is a finite type system, which is not solvable for the parameter $g_{0}$ and hence not solvable for a generic metric $g$.
@@ 5. Discussion: on the selection principle
Conformal geodesics extend naturally to (and through) conformal infinity. Their behavior at the conformal infinity uncovers the structure of asymptotically simple solutions of the Einstein field equations [22, 11]. In infinity this family of curves carries only a projective family of parametrizations, which makes un-parametrized conformal geodesics at the three-dimensional spatial boundary physically meaningful.
Our main result states that the variationality of conformal geodesics is the selection principle for physical dimension three. In all other dimensions (including dimension two and parametrized version in dimension three) the equation is not variational. Here we test only the first variation of conformal geodesics. Indeed, even in the flat case, there exist infinitely many conformal geodesics through any two (sufficiently close) points and thus the second variation does not seem to be relevant.
@@
[…]
For instance, in the case of conformal geodesic equation $\mathcal{E}$ a nondegenerate conformal invariant Lagrangian of order 2 was proposed in [3], which up to divergence coincides with the one investigated further in [8]. It was noted in [3] that its Euler-Lagrange equation is of forth order; this equation $\mathcal{E}^{\triangledown}\supset\mathcal{E}$ containing the equation of conformal geodesics was later studied in [23, 20]. An example of smaller equation $\mathcal{E}_{\vartriangle}$ (for Kähler metric $g$) is given by Kähler magnetic geodesics discussed in [8].
@@
Let the equation be given by $\mathcal{E}=\{F_{i}=0\}_{i=1}^{m}$, where $F_{i}$ are (nonlinear) differential operators on $u=(u^{j})_{j=1}^{n}$ (usually, for determined systems, $m=n$). Then $L^{\triangledown}=\sum_{i=1}^{m}v^{i}F_{i}$ for additional dependent variable $v=(v^{i})_{i=1}^{m}$. Indeed, the equations $\delta_{v^{i}}L^{\triangledown}=F_{i}=0$ for $1\leq i\leq n$ imply $\mathcal{E}$. The other Euler-Lagrange equations $\delta_{u^{j}}L^{\triangledown}=\sum_{i=1}^{n}v^{i}\delta_{u^{j}}F_{i}=0$ should be considered as constraints on $v$. Taken together these two sets of equations define what is known as cotangent covering $\mathcal{T}^{*}\mathcal{E}$ over $\mathcal{E}$, cf. [25]. Its linearization is given by the operator $\ell_{\mathcal{E}}\oplus\ell_{\mathcal{E}}^{*}$ that is self-adjoint and hence (see e.g. [1, 24]) $\mathcal{T}^{*}\mathcal{E}$ is a variational equation. ∎
Similarly, any differential equation $\mathcal{E}$ (with more than one-dimensional solution space) contains a reduction to scalar second order ODE (path geometry) that is variational (a result by Sonin of 1886, see [12]). In other words, under mild conditions, every equation $\mathcal{E}$ contains a variational intermediate integral $\mathcal{E}_{\vartriangle}$.
Our final remark concerns variations with prescribed boundary conditions. With the standard approach one usually assumes fixed endpoint and derivatives of variations at it. Non-standard variational principle for conformal geodesics in [8] restricts the variations $V$ along the curve by $V^{\prime}(t_{0})=0$, $V^{\prime}(t_{1})=0$ as well as the boundary condition $B(V)|_{t_{0}}^{t_{1}}=-K(V)$, where $B$ is a certain second order differential operator. For the flat metric $K=0$. Consider, for instance, a one-parametric family of conformal circles in ${\mathbb{R}}^{n}$, with center $(0,\tfrac{1-c^{2}}{2c},0,\dots,0)$ and radius $\tfrac{c^{2}+1}{2c}$, passing through the points $(\pm 1,0,\dots,0)$ for $t=\pm 1$:
@@
$\bigl.B(V)\bigr|_{t=-1}^{t=+1}=\frac{c^{2}-1}{c^{2}+1}\bigl(V_{1}^{\prime\prime}(1)-V_{1}^{\prime\prime}(-1)\bigr)-\frac{2c}{c^{2}+1}\bigl(V_{2}^{\prime\prime}(1)+V_{2}^{\prime\prime}(-1)\bigr)-\frac{2}{c^{2}+1}\bigl(V_{1}(1)-V_{1}(-1)\bigr).$
@@
Let us finish with a historical remark. The notion of conformal geodesics has appeared in literature in at least three different contexts. The first corresponds to the title of Fialkov’s paper [10], where he mentions the preceeding work by Schouten of 1928, and there conformal geodesics are defined as a family of curves, more precisely a path structure that is metrizable within the given conformal class. No particular curve from this family can be characterized by a relation to the conformal structure. Indeed, it was shown in Lemma 2 of [9] that any curve is locally a geodesic for a metric in a given conformal class.
Next, conformal geodesics were introduced as extremals for the so-called conformal arclength functional by Musso in [19] and then extended in [17] (this action corresponds to the simplest differential invariant one-form). These curves are variational by construction, but their Euler-Lagrange equation is of order six.
Finally, our usage of conformal geodesics is related to [11], which is based on the work by Schmidt of 1986 with a reference to the preceeding paper by Ogiue of 1967; it was re-used in physical literature by Tod in 2012 and later, see references in [8]. In mathematical literature this equation was introduced in [26] in 1957; later in [3] it was more conservatively called the equation of conformal circles. Given the main result of this paper it may be questioned which nomenclarure is more justified.
Acknowledgment. I would like to thank Vladimir Matveev, Wijnand Steneker, Eugene Ferapontov, Iosif Krasilshchik, Konstantin Druzhkov, Michael Eastwood and Josef Silhan for useful discussions. The research was supported by the Tromsø Research Foundation (project “Pure Mathematics in Norway”) and the UiT Aurora project MASCOT.